PLEASE HELP The rectangular form for the posar equation r = -2csc(0 with a line through it) is
a. y = x-2
b. y = 2
c. y = 2x
d. y = -2
thank you
: PLEASE HELP The rectangular form for the posar equation r = -2csc(0 with a line through it) is
a. y = x-2
b. y = 2
c. y = 2x
d. y = -2
thank you

Answer by stanbon(19007)

The rectangular form for the posar equation r = -2csc(0 with a line through it) is
a. y = x-2
b. y = 2
c. y = 2x
d. y = -2
-----------------
Conversion formulas:
r = sqrt(x^2+y^2)
theta = tan^-1(y/x)
------------------------------
Note: If tangent is y/x then csc is r/y
------------------------------
Substituting you get:
r =-2(r/y
Divide both sides by r to get:
1 = -2/y
y = -2
================
Cheers,
Stan H.